# Machine Learning needs its own Randomness Standard: Randomised Smoothing and PRNG-based attacks

### v1

@@ 5.3 Bit-flipping PRNG attacker
In this subsection we evaluate the three attacks described in Section 4.3: negative-kurtosis, skewness and positive-kurtosis. The parameter values chosen for evaluation are $\alpha\in\left\{1,2,3,4\right\}$, $\beta\in\left\{0,1,2,4\right\}$ and $\gamma\in\left\{0,1,2,4\right\}$ for each of the attacks respectively. Note that neither attack can be detected by a naive defender. Appendix D shows the relative accuracy results, that do not deviate significantly from the baseline. Relative certification radii for each attack are shown in Table 2, with the skewness attack achieving the highest manipulated radii, followed by the positive-kurtosis. The negative-kurtosis attack performed worst, significantly altering the radius for only 4% of images, even with $\alpha=1$.
@@ 5.4 Defences
NIST testsTable 3 reports the number of bitstreams that passed each test out of 1000, where each bitstream was of length $10^{6}$. This is excluding the random excursions and random excursions variant test, for which the NIST utility decides the number of bit streams to consider based on previous test results. The recommendation from NIST is that a good PRNG should pass at least 980 out of 1000 runs of each test. For the random excursion tests, the pass rate is reported based on the number of tests run. In addition to the attacked PRNGs, the results are also reported for the four most popular PRNGs currently in use: MT19937, PCG64, Philox and SFC64. While the other three manage to pass all tests, the MT19937 PRNG fails most of them, which is good news, since MT19937 is not recommended for any use case where a secure PRNG is required.
Focusing on the results for $\beta=4$ and $\gamma=4$, the tests that were able to successfully detect both the skewness and positive-kurtosis attacks with high confidence are the frequency test, cumulative sums test and the runs test. It must be noted that the minimum recommended input size by NIST for these three tests is 100, which is significantly lower than the input size of $10^{6}$ used for generating the results in Table 3. Figure 3 shows the pass rate for these tests for different sample sizes ranging from $10^{2}$ to $10^{6}$ for the attacked PRNGs with $\beta=4$ and $\gamma=4$. Results show a significant drop in pass rate only after the input size is increased above $10^{5}$, demonstrating a clear need to re-evaluate the recommended parameters and input sizes suggested by NIST for safety-critical ML.
Another important result is that for the negative-kurtosis attack with $\alpha=3$. It was able to get a significantly higher pass rate compared to the rest of the PRNGs for the same attack. It goes to show that statistical tests cannot always be relied upon and bad PRNGs sometimes get high pass rates.
@@
Figure 3: These figures show the NIST test pass rates for the frequency, cumulative sums and runs tests, the three tests that managed to detect all the attacked PRNGs with high confidence at an input size of $10^{6}$. These plots show how the pass rates vary as the input size is increased from the minimum recommended value by NIST, $10^{2}$, to $10^{6}$ for the two attacks that were the hardest to detect: skewness attack with $\beta=4$ and positive-kurtosis attack with $\gamma=4$. The attacks become detectable only after the input size is increased to $>10^{5}$.
@@
Shapiro test The results from the NIST tests demonstrate that it is difficult to detect attacks against PRNGs that alter the distribution once it is transformed to the normal distribution by using existing tests developed for ensuring that PRNGs are cryptographically secure. While attacks such as the negative kurtosis attack that significantly alter the distribution of raw random integers produced by the PRNG can be detected, more sophisticated attacks such as the skewness attack, and the negative kurtosis attack escape detection since they modify specific bits that have a greater impact on the normal distribution transformation. Therefore, the next class of defences explored here focuses on detecting the attack after the transformation and is only performed for the skewness and positive kurtosis attacks.
@@ 6 Discussion
This work successfully demonstrates the feasibility of an attack on the pseudo-random number generator in an ML library to manipulate the certified radius obtained for the industry standard Randomised Smoothing technique. By altering the bit stream produced by the random number generator, kurtosis or skewness can be introduced into the noise distribution used. This can significantly affect the certified radius and thus make the end user over-estimate the real robustness. This work adds to a small but growing area of research into machine learning security that looks beyond the attack surfaces of the models themselves.
Technical stack ML security is just like traditional security, in that practitioners need to consider the whole technical stack and not just limit themselves to the ‘ML’ piece of data collection, training and inference-time attacks. An ML model runs on the same kind of hardware and software layers as every other piece of software and is therefore just as vulnerable to these ‘traditional’ attack surfaces. The literature already notes that attacks can come from ML compilers [11], underlying platforms [53, 8], model architectures [7], or even quantisation granularity [43]. Yet there is no standard guide for secure ML deployment.
@@
On passing tests Bassham et al. stress [5] that by the very nature of statistical tests, some good PRNGs will fail them, while some bad PRNGs will pass. Therefore, the design of a new PRNG should be mathematically and architecturally sound, allow for external scrutiny, and its implementation must be studied carefully for backdoors. Other applications of randomness in ML are also likely to have specific vulnerabilities and will need explicit tests designed to target them e.g. in differential privacy.
@@
New Standards We argue that ML needs its own standard for randomness. This must be informed by a better threat model, as well as application domain knowledge. For example, if a given ML application uses the PRNG to sample Gaussian noise, then this property must be explicitly verified. Appendix F shows that each normality test detects different attacks, while no test detects all of the attacks reliably. A comprehensive set of tests should be mandated and should become a part of the standard.
@@
ML brings up novel issues that were never a concern of cryptographers. In particular, we need to revisit issues around floating-point representations of random numbers – Mironov noted these in 2011 [49], and Zhuang et al. found that similar attacks work in 2022 [70]. Our attacks further demonstrate the importance of these issues. In cryptographic applications, key generation is performed relatively rarely, so an application developer who is suspicious of the quality of random numbers supplied by the platform will pass them through a PRNG of their own construction, which typically maintaine a pool of randomness and used one-way functions both to update this pool and to draw key material from it. However this uses several Kb of memory and several thousand instructions for every pseudorandom value drawn. Many ML applications cannot afford to do this as they make intensive use of computation, draw many random values, and can waste neither compute cycles nor GPU-adjacent RAM. This rules out the use of application-specific PRNGs, and also rules out continuous testing of the random numbers supplied by the platform – running the NIST suite takes hours.
@@
As such testing is impractical at runtime we may well therefore need separate test suites for system certification and to detect runtime attacks, as in some cryptographic systems; but that may not be enough. In critical applications, designers may need to understand that mechanisms for shaping bitstream statistics are within the trusted computing base. System evaluation will also have to take account of this, and demand assurance against the kind of attacks discussed here.
@@ 7 Conclusion
Machine learning systems rely on randomness for many purposes, ranging from differential privacy, through selecting representative subsets of data for training, to privacy amplification in federated learning and the generation of synthetic data. Yet the consequences of poor random number generators have remained unstudied; common tools optimise random number generation for speed rather than security, and no attention is paid to the possibility that adversaries might manipulate random number sources to undermine model safety or security. In this paper we presented two proof-of-concept attacks on randomised smoothing, showing first that replacing a Gaussian random generator with a Laplacian one is effective if detectable, and second that flipping targeted bits in a Gaussian random generator is also effective but much more covert, in the sense of being very hard to detect using standard NIST randomness tests with default parameters.
The consequences, and future directions of research, are three-fold. First of all, there is a real need to define what security means for an ML PRNG, just as there has been extensive study of the requirements for cryptographic PRNGs. Second, we need to put more effort into exploring attacks that modify the random bit-stream in less predictable ways to improve attack performance and reduce the probability of detection. Finally, we need proper standards for randomness in ML systems, and we believe that a careful study of both attack and defence is the necessary foundation.
@@
Appendix A Sampling from Normal Distribution
Figure 5: Rectangle-wedge-tail method
Appendix B Implementation of Bit-flipping PRNG attacks
@@
Appendix C NIST Test Suite
  1. 1.
    Frequency (Monobit) Test: The purpose of this test of to verify whether the proportion of 1s and 0s in the bit stream is as expected from a random bit stream. The input size recommendation is at least 100 bits.
  2. 2.
    Frequency Test within a Block: This test validates that the number of ones in every M-bit block is roughly $M/2$. This is equivalent to the previous test for M=1. The input size recommendation is at least 100 bits, with $M>20$, such that $M\geq 0.1n$.
  3. 3.
    Runs Test: A run is defined as a sequence of identical bits, bounded by the opposite bit, before and after. This test validates that the number of runs of 1s and 0s of every length is as expected. The input size recommendation is at least 100 bits.
  4. 4.
    Longest Run of Ones in a Block: This test checks for the longest run of 1s in every M-bit block. The test has three possible values of $M$, and the corresponding minimum input size recommendation is specified below:
    Minimum $n$ $M$
    128 8
    6272 128
    750,000 1,000
  5. 5.
    Binary Matrix Rank Test: This test looks for linear dependence between equal-sized substrings of the bit stream. The test is hard coded for $32\times 32$ square matrices, and the recommendation is that there be at least 38 matrices in the bit stream, which means $n$ must be at least 38,912.
  6. 6.
    Discrete Fourier Transform Test: This test attempts to detect patterns in the Fourier transform of the bit stream, in the frequency space. The minimum recommended input size is 1000.
  7. 7.
    Non-overlapping Template Mapping Test: This test uses pre-determined $m$-bit sequences to check if any pattern appears more often than expected in the bit stream. Templates up to $m=10$ are included in the test suite, and $m\geq 9$ is recommended to get meaningful results. This test was run for $m=9$, for bit streams of length $10,000$. Since there are $148$ $9$-bit templates, and the test is run once for each template, and the minimum pass rate out of all templates is reported.
  8. 8.
    Overlapping Template Matching Test: This is the same test as the previous one, except it looks for overlapping matches for substrings as well. The minimum input size recommendation is $10^{6}$. As with the previous test, the minimum pass rate is reported.
  9. 9.
    Maurer’s Universal Statistic Test: This test checks the extent to which the bit stream can be losslessly compressed. The more a sequence can be compressed, the less random it is. The minimum recommended input size is 387,840.
  10. 10.
    Linear Complexity Test: This test determines the length of a linear feedback shift register, and verifies if the sequence is complex enough to be random. The minimum recommended input size is $10^{6}$.
  11. 11.
    Serial Test: This test determines whether the occurrence of all $m$-bit patterns is equally likely in the bit stream. For $m=1$, this is the same as the frequency test. The recommendation is that $m$ and $n$ be chosen such that $m<\log_{2}n-2$.
  12. 12.
    Approximate Entropy Test: This test compares the occurrences of overlapping blocks of sizes $m$ and $m+1$. The recommendation is to choose $m$ and $n$ such that $m<\log_{2}n-5$.
  13. 13.
    Cumulative Sums Test: This test considers a random walk such that 0 is -1, an 1 is +1, and determines if such random walks determined by partial sequences in the bit stream deviate significantly from zero. The input size recommendation is a minimum of 100 bits.
  14. 14.
    Random Excursions Test: This test determines the number of random walk sequences that visit a specific state, and whether that deviates from random behaviour. 8 states are considered, -4, -3, -2, -1, +1, +2, +3 and +4. The input size recommendation for this test is at least $10^{6}$ bits. Eight different results for each state are reported, and the minimum pass rate is considered.
  15. 15.
    Random Excursions Variant Test: This is the same as the previous test, but for 18 different states, from -9 to +9. The minimum recommended input size is at least $10^{6}$ bits.
Appendix D Relative Accuracy for Bit-flipping PRNG Attacks
@@
Appendix E Baseline QQ plot for PCG64 PRNG
@@
Appendix F Other normality tests
Tables 8, 9, 10, 11, 12 and 12 show the results of the following tests for normality with different sample sizes on the PRNG attacks:
  1. 1.
    D’Agostino’s K-squared test
  2. 2.
    Kolmogorov-Smirnov test
  3. 3.
    Cramer-von Mises criterion
  4. 4.
    Jarque-Bera test
  5. 5.
    Shapiro-Wilk test
  6. 6.
    Lilliofors test
@@ Appendix F Other normality tests
The results are reported in a similar manner to the NIST test results in Table 3. 1000 instances of each test were run for every choice of sample size from $10^{2}$ to $10^{6}$. A test was passed if the $p$-value reported was greater than 0.01. Setting the same threshold as that recommended by NIST, with 980/1000 as the minimum pass rate, it is clear that different tests perform differently depending on the type of attack. The negative kurtosis attack, which was the easiest to detect with the NIST tests, is now the hardest as it alters the distribution the least. In order to get conclusive results with low pass rates for all PRNGs, the sample size had to be set to at least $10^{5}$ in Table 11. This highlights the importance of using both types of testing: on the random bit stream and after transformation, as different types of attacks may be detectable.
D’Agostino Kolmogorov-Smirnov Cramer-von Mises Jarque-Bera Shapiro Lilliefors
baseline 976 988 990 982 990 993
$\alpha=1$ 982 975 985 994 993 981
$\alpha=2$ 994 987 987 995 996 996
$\alpha=3$ 986 988 989 991 991 993
$\alpha=4$ 983 987 986 985 990 996
$\beta=0$ 225 0 0 207 8 318
$\beta=1$ 829 0 0 809 778 631
$\beta=2$ 939 262 281 931 931 875
$\beta=4$ 972 825 833 970 968 959
$\gamma=0$ 715 867 819 793 803 983
$\gamma=1$ 885 972 988 856 916 982
$\gamma=2$ 933 989 991 912 952 989
$\gamma=4$ 962 993 993 949 975 990
Table 8: Test results for sample size = 100. Smaller number are better for the defender – 1000 means no attack is detected, 0 means all of the attacks are detected. Highlighted are the results that according to NIST guidelines pass the randomness tests with suggested threshold of 980.
D’Agostino Kolomogorov-Smirnov Cramer-von Mises Jarque-Bera Shapiro Lilliefors
baseline 988 992 995 983 989 994
$\alpha=1$ 887 749 792 968 894 855
$\alpha=2$ 981 964 977 997 986 978
$\alpha=3$ 982 980 986 993 990 991
$\alpha=4$ 986 987 988 993 987 986
$\beta=0$ 0 0 0 0 0 0
$\beta=1$ 12 0 0 5 0 0
$\beta=2$ 257 0 0 248 41 0
$\beta=4$ 726 0 0 718 566 276
$\gamma=0$ 251 0 0 260 0 857
$\gamma=1$ 400 80 35 298 411 986
$\gamma=2$ 695 715 699 590 733 986
$\gamma=4$ 903 957 966 853 929 994
Table 9: Test results for sample size = 1000. Smaller number are better for the defender – 1000 means no attack isdetected, 0 means all of the attacks are detected. Highlighted are the results that according to NIST guidelines pass therandomness tests with suggested threshold of 980.
D’Agostino Kolomogorov-Smirnov Cramer-von Mises Jarque-Bera Lilliefors
baseline 989 984 990 989 985
$\alpha=1$ 15 0 0 21 0
$\alpha=2$ 713 274 205 762 460
$\alpha=3$ 932 747 792 953 819
$\alpha=4$ 960 884 932 973 916
$\beta=0$ 0 0 0 0 0
$\beta=1$ 0 0 0 0 0
$\beta=2$ 0 0 0 0 0
$\beta=4$ 0 0 0 0 0
$\gamma=0$ 1 0 0 1 0
$\gamma=1$ 0 0 0 0 184
$\gamma=2$ 1 0 0 1 502
$\gamma=4$ 78 0 0 56 843
Table 10: Test results for sample size = 10000. Smaller number are better for the defender – 1000 means no attack isdetected, 0 means all of the attacks are detected. Highlighted are the results that according to NIST guidelines pass therandomness tests with suggested threshold of 980.
D’Agostino Kolomogorov-Smirnov Cramer-von Mises Jarque-Bera Lilliefors
baseline 991 992 993 992 977
$\alpha=1$ 0 0 0 0 0
$\alpha=2$ 0 0 0 0 0
$\alpha=3$ 22 0 0 25 0
$\alpha=4$ 291 0 0 314 30
$\beta=0$ 0 0 0 0 0
$\beta=1$ 0 0 0 0 0
$\beta=2$ 0 0 0 0 0
$\beta=4$ 0 0 0 0 0
$\gamma=0$ 0 0 0 0 0
$\gamma=1$ 0 0 0 0 0
$\gamma=2$ 0 0 0 0 0
$\gamma=4$ 0 0 0 0 0
Table 11: Test results for sample size = 100000. Smaller number are better for the defender – 1000 means no attack isdetected, 0 means all of the attacks are detected. Highlighted are the results that according to NIST guidelines pass therandomness tests with suggested threshold of 980.
D’Agostino Kolomogorov-Smirnov Cramer-von Mises Jarque-Bera Lilliefors
baseline 996 984 983 996 946
$\alpha=1$ 0 0 0 0 0
$\alpha=2$ 0 0 0 0 0
$\alpha=3$ 0 0 0 0 0
$\alpha=4$ 0 0 0 0 0
$\beta=0$ 0 0 0 0 0
$\beta=1$ 0 0 0 0 0
$\beta=2$ 0 0 0 0 0
$\beta=4$ 0 0 0 0 0
$\gamma=0$ 0 0 0 0 0
$\gamma=1$ 0 0 0 0 0
$\gamma=2$ 0 0 0 0 0
$\gamma=4$ 0 0 0 0 0
Table 12: Test results for sample size = 1000000. Smaller number are better for the defender – 1000 means no attack isdetected, 0 means all of the attacks are detected. Highlighted are the results that according to NIST guidelines pass therandomness tests with suggested threshold of 980.
CyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 5.3 Bit-flipping PRNG attacker
In this subsection we evaluate the three attacks described in Section 4.3: negative-kurtosis, skewness and positive-kurtosis. The parameter values chosen for evaluation are $\alpha\in\left\{1,2,3,4\right\}$, $\beta\in\left\{0,1,2,4\right\}$ and $\gamma\in\left\{0,1,2,4\right\}$ for each of the attacks respectively. Note that none of these attacks can be detected by a naive defender. Appendix B shows the relative accuracy results, that do not deviate significantly from the baseline. Relative certification radii for each attack are shown in Table 2, with the skewness attack achieving the highest manipulated radii, followed by the positive-kurtosis. The negative-kurtosis attack performed worst, significantly altering the radius for only 4% of images, even with $\alpha=1$.
@@ 5.4 Defences
NIST testsTable 3 reports the number of bitstreams that passed each test out of 1000, where every bitstream was of length $10^{6}$. This excludes the random excursions and random excursions variant test, for which the NIST utility decides the number of bit streams to consider based on previous test results. The recommendation from NIST is that a good PRNG should pass at least 980 out of 1000 runs of each test. For the random excursion tests, the pass rate is reported based on the number of tests run. In addition to the attacked PRNGs, the results are also reported for the four most popular PRNGs currently in use: MT19937, PCG64, Philox and SFC64. While the other three manage to pass all tests, the MT19937 PRNG fails most of them, which is good news, since MT19937 is not recommended for any use case where security is required.
Focusing on the results for $\beta=4$ and $\gamma=4$, the tests that were able to successfully detect both the skewness and positive-kurtosis attacks with high confidence are the frequency test, cumulative sums test and the runs test. It must be noted that the minimum recommended input size by NIST for these three tests is 100, which is significantly lower than the input size of $10^{6}$ used for generating the results in Table 3. Figure 4 shows the pass rate for these tests for different sample sizes ranging from $10^{2}$ to $10^{6}$ for the attacked PRNGs with $\beta=4$ and $\gamma=4$. Results show a significant drop in pass rate only after the input size is increased above $10^{5}$, demonstrating a clear need to re-evaluate the recommended parameters and input sizes suggested by NIST for safety-critical applications.
Another important result is that for the negative-kurtosis attack with $\alpha=3$. It was able to get a significantly higher pass rate compared to the rest of the PRNGs for the same attack. It goes to show that statistical tests cannot always be relied upon and bad PRNGs sometimes get a high pass rates.
@@
Figure 4: These figures show the NIST test pass rates for the frequency, cumulative sums and runs tests, the three tests that managed to detect all the attacked PRNGs with high confidence at an input size of $10^{6}$. These plots show how the pass rates vary as the input size is increased from the minimum recommended value by NIST, $10^{2}$, to $10^{6}$ for the two attacks that were the hardest to detect: the skewness attack with $\beta=4$ and the positive-kurtosis attack with $\gamma=4$. The attacks become detectable only after the input size is increased to $>10^{5}$.
@@
Shapiro test The results from the NIST tests demonstrate that it is difficult to detect attacks against PRNGs that alter the distribution once it is transformed to the normal distribution by using existing tests developed for ensuring that PRNGs are cryptographically secure. While attacks such as the negative kurtosis attack that significantly alter the distribution of raw random integers produced by the PRNG can be detected, more sophisticated attacks such as the skewness attack and the negative kurtosis attack escape detection since they modify bits that have a greater impact on the normal distribution transformation. Therefore, the next class of defences explored here focuses on detecting the attack post-transformation, and is only performed for the skewness and positive kurtosis attacks.
@@ 6 Discussion
This work successfully demonstrates the feasibility of an attack on the pseudo-random number generator in an ML library to manipulate the certified radius obtained for the industry standard Randomised Smoothing technique. By altering the bit stream produced by the random number generator, kurtosis or skewness can be introduced into the noise distribution used. This can significantly affect the certified radius and thus make the end user over-estimate the real robustness. This work adds to a small but growing area of research into machine learning security that looks beyond the attack surfaces of the models themselves and considers systems that enable them.
Broader Impact ML security is just like traditional security, in that practitioners need to consider the whole technical stack and not just limit themselves to the ‘ML’ parts of data collection, training and inference-time attacks. An ML model runs on the same kind of hardware and software layers as every other piece of software and is therefore just as vulnerable to these ‘traditional’ attack surfaces. The literature already notes that attacks can come from ML compilers [10], underlying platforms [51, 8], model architectures [7], or even quantisation granularity [43]. Yet there is no standard guide for secure ML deployment, making responsible deployment challenging.
@@
Limitations and Realism Although access to randomness generators is a strong adversary assumption, it is not unrealistic. The dual EC generator was backdoored by NSA in the past to break cryptography that used it. In this paper we demonstrate that randomness can similarly be used to break ML applications and it should be included in threat models.
Practical Attacks In this paper we launch our attacks using the random number generator – in practice they can also be utilised at different points in the technical stack. For example, an adversary can use BitFlips to change random numbers passing over the PCI-e bus [51]; if randomness comes from the host platform, they could use VPGA attacks [8] or Rowhammer [30] to achieve the same effect.
@@
On passing tests Bassham et al. stress [5] that by the very nature of statistical tests, some good PRNGs will fail them, while some bad PRNGs will pass. For example, we find that cryptographically secure PCG64 fails Lilliefors test at $10^{5}$ samples (Table 7). Therefore, the design of a new PRNG should be mathematically and architecturally sound, allow for external scrutiny, and its implementation must be studied carefully for backdoors. Other applications of randomness in ML are also likely to have specific vulnerabilities and will need tests designed to target them e.g. in differential privacy.
@@
Better Standards We argue that ML needs better standards for randomness. This must be informed by a better threat model, as well as application domain knowledge. While a cryptographically secure PRNG can sometimes meet this requirement, ML developers often look for better performance, and their requirements for randomness differ in other ways from those of cryptographers. ML requires that the sampling distribution is accurate whereas cryptographers aim to reduce predictability. So while there may be an overlap between the requirements for the two applications, there is still a need for ML-specific randomness guidance. For example, if a given ML application uses the PRNG to sample Gaussian noise, then this property must be explicitly verified. Appendix D shows that each normality test detects different attacks, while no test detects all of them reliably. A comprehensive test suite should become part of the standard with a well-calibrated sample size e.g. $10^{5}$ for Gaussian noise-dependent ML workloads and multiple Guassian noise-specific tests.
@@
ML brings up novel issues that were never a concern of cryptographers. In particular, we need to revisit issues around floating-point representations of random numbers – Mironov noted these in 2012 [47], and Zhuang et al. found that similar attacks work in 2022 [69]. Our attacks further demonstrate the importance of these issues. In cryptographic applications, key generation is performed relatively rarely, so an application developer who is suspicious of the quality of random numbers supplied by the platform will pass them through a PRNG of their own construction, which typically maintains a pool of randomness and used one-way functions both to update this pool and to draw key material from it. However, this uses several Kb of memory and several thousand instructions for every pseudorandom value drawn. Many ML applications cannot afford to do this as they make intensive use of computation, draw many random values, and can waste neither compute cycles nor GPU-adjacent RAM. This rules out the use of application-specific PRNGs, and also rules out continuous testing of the random numbers supplied by the platform – running the NIST suite already takes hours.
@@
As such testing is impractical at runtime we may well therefore need separate test suites for system certification and to detect runtime attacks, as in some cryptographic systems; but that may not be enough. In critical applications, designers may need to understand that mechanisms for shaping bitstream statistics are within the trusted computing base. System evaluation will also have to take account of this and demand assurance against the kind of attacks discussed here.
@@ 7 Conclusion
Machine learning systems rely on randomness for many purposes, ranging from differential privacy, through selecting representative subsets of data for training, to privacy amplification in federated learning and the generation of synthetic data. Yet the consequences of poor random number generators have remained unstudied. Common tools optimise random number generation for speed rather than security, and no attention is paid to the possibility that adversaries might manipulate random number sources to undermine model safety or security. In this paper we presented two proof-of-concept attacks on randomised smoothing, showing first that replacing a Gaussian random generator with a Laplacian one is effective if detectable, and second that flipping targeted bits in a Gaussian random generator is also effective but much more covert, in the sense of being very hard to detect using standard NIST randomness tests with default parameters.
The consequences, and future directions of research, are three-fold. First of all, there is a real need to define what security means for an ML PRNG, just as there has been extensive study of the requirements for cryptographic PRNGs. Second, we need to put more effort into exploring attacks that modify the random bit-stream in less predictable ways to improve attack performance and reduce the probability of detection. Finally, we need proper standards for randomness in ML systems and indeed in statistics generally – for which a careful study of both attack and defence is the necessary foundation.
@@
Appendix A Implementation of Bit-flipping PRNG attacks
@@
Appendix B Relative Accuracy for Bit-flipping PRNG Attacks
@@
D’Agostino Kolmogorov-Smirnov Cramer-von Mises Jarque-Bera Shapiro Lilliefors PCG64 - - - - - $10^{5}$ $\alpha=1$ $10^{3}$ $10^{2}$ $10^{3}$ $10^{3}$ $10^{3}$ $10^{3}$ $\alpha=2$ $10^{4}$ $10^{3}$ $10^{3}$ $10^{4}$ $10^{4}$ $10^{4}$ $\alpha=3$ $10^{4}$ $10^{3}$ $10^{4}$ $10^{4}$ $10^{4}$ $10^{4}$ $\alpha=4$ $10^{4}$ $10^{4}$ $10^{4}$ $10^{4}$ $10^{4}$ $10^{4}$ $\beta=0$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $\beta=1$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $\beta=2$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $\beta=4$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $\gamma=0$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{2}$ $10^{4}$ $\gamma=1$ $10^{2}$ $10^{2}$ $10^{3}$ $10^{2}$ $10^{2}$ $10^{4}$ $\gamma=2$ $10^{2}$ $10^{3}$ $10^{3}$ $10^{2}$ $10^{2}$ $10^{4}$ $\gamma=4$ $10^{2}$ $10^{3}$ $10^{3}$ $10^{2}$ $10^{2}$ $10^{4}$
Table 7: This table reports the minimum sample size for which tests started to fail for each PRNG compared to PCG64. 1000 tests were run for increasing sample sizes from $10^{2}$ to $10^{6}$ following the same pass criteria as the NIST test suite in Table 3. A sample size of $\mathbf{10^{5}}$ was required to see every attacked PRNG conclusively fail all 6 tests.
@@
Appendix C Baseline QQ plot for PCG64 PRNG
@@
Appendix D Other normality tests
@@
Table 7 show the results of the following tests for normality with different sample sizes on the PRNG attacks: D’Agostino’s K-squared test, Kolmogorov-Smirnov test, Cramer-von Mises criterion, Jarque-Bera test, Shapiro-Wilk test, and Lilliofors test. 1000 instances of each test were run for varying sample sizes from $10^{2}$ to $10^{6}$. A test was passed if the $p$-value reported was greater than 0.01 and the overall pass threshold was set as 980/1000 tests according to NIST’s recommendations. Table 7 reports the minimum sample size for which each test failed for the different attacked PRNGs. Test results for a baseline PCG64 PRNG are also reported. It is clear that tests perform differently depending on the type of attack. The negative kurtosis attack, which was the easiest to detect with the NIST tests, is now the hardest as it alters the distribution the least. In order to get conclusive results with low pass rates for all PRNGs, the sample size had to be set to at least $10^{5}$. This highlights the importance of using both types of testing: on the random bit stream and after transformation, as different types of attacks may be detectable.